OpenStudy (anonymous):

find the derivative of sin^-1(2x+3)

OpenStudy (anonymous):

what i did was change Sin^-1(2x+3) to Sin^-1(x+3/1/2) so its in the form of x/a

OpenStudy (anonymous):

\[y = \sin^{-1}(2x+3) \iff \sin(y) = 2x+3 \iff \cos(y)\frac{dy}{dx} = 2 \] \[\iff \frac{dy}{dx} = \frac{2}{\cos(\sin^{-1}(2x+3))} = \frac{2}{\sqrt{1-(2x+3)^2}}\] i think <.< never liked calculus.

OpenStudy (anonymous):

im probably doing it the long way. most likely theres some short formula out there.

OpenStudy (anonymous):

well there this formula Sin^-1(x/a) = 1/sqrt(a^2-x^2)

OpenStudy (anonymous):

use the chain rule then. let: \[\frac{d}{dx}\sin^{-1}\left(\frac{f(x)}{a}\right) = \frac{1}{\sqrt{a^2-f(x)^2}}*f'(x)\]

OpenStudy (anonymous):

f(x) = 2x+3, a = 1

OpenStudy (anonymous):

my answer ended up being : 2/sqrt(1-4(x+3)^2)

OpenStudy (anonymous):

but i think the x has to be isolated doesn't it?

OpenStudy (anonymous):

using the chain rule, you want to keep the function together.

OpenStudy (anonymous):

whatever you define the function to be. Since im defining it as f(x) = 2x+3, i dont want to separate it.

OpenStudy (anonymous):

ok i see thanks